# Composing Concepts from Images and Videos via Concept-prompt Binding

### v1

@@ 4.2.1 Quantitative Comparisons
We construct 40 test cases with images and videos from the DAVIS [davis] dataset and the Internet for evaluation. Both automatic metrics and human evaluations are adopted for assessing the concept composition performance. For automatic metrics, we use CLIP-T to measure the alignment between the generated video and the textual prompt with CLIP [clip] feature similarities, and choose DINO-I to quantify the preservation of visual concepts with the harmonic mean of DINO [dino] feature similarities between the composed video and all visual inputs. For human evaluations, we asked 28 volunteers to rate the composed video in the following aspects with a 5-point Likert scale: 1) Concept Preservation: how well the composed video preserves the concepts from corresponding visual sources. 2) Prompt Fidelity: how well the composed video follows the input prompt. 3) Motion Quality: the motion quality of the composed video considering motion smoothness, consistency, naturalness, etc. We compute the average score of the 3 aspects as the Overall Quality. Please refer to the supplementary for more details on user study settings.
Table 1: Quantitative Comparisons with Prior Arts4.2.1). Results in bold are the best. † Implemented on Wan2.1 [wan].
@@
As displayed in Tab. 1, BiCo consistently outperforms all other methods in both automatic metrics and human evaluations. Compared to the prior art DualReal [wang2025dualrealadaptivejointtraining], our method achieves a +54.67% improvement on the subjective Overall Quality. In addition, BiCo also supports the extraction of non-object concepts, learning multiple concepts from a single input, arbitrary image/video input types, and flexible concept composition via prompt manipulation, where previous methods fall short.
@@ 4.2.2 Qualitative Comparisons
We visualize the composed videos in Fig. 5 to provide an intuitive comparison with other methods. It shows a creative motion transfer task, where Textual Inversion [textualinversion] and DreamVideo [dreamvideo] fails to combine the visual concepts. DualReal [wang2025dualrealadaptivejointtraining] does not accurately follow the prompt to compose the concepts, and the generated video is almost static. Although DB-LoRA [dreambooth] mostly follows the designated prompt to integrate visual concepts, there are significant drifts of visual concepts from the original inputs. BiCo best composes the visual concepts according to the given prompt while maintaining the visual concept consistency with the input image and video.
@@ 5 Conclusion and Discussion
[…]
@@
  • §C further ablates the two-stage inverted training strategy.
  • §D illustrates more qualitative comparisons with previous methods.
  • §E performs another case study to facilitate the understanding of different components of BiCo.
  • §F further discusses the limitations with failure cases and the societal impacts of BiCo.
Please refer to the supplemental page for video results.
@@ Appendix A Detailed Prompts for DAM
In the prompt diversification process, we utilize a powerful VLM Qwen2.5-VL [bai2025qwen25vltechnicalreport] to generate diversified concept prompts while retaining the key conceptual words unchanged. During the key concept extraction stage, the VLM is asked to extract essential spatial and temporal concepts from the visual inputs. For image inputs, we use the following textual prompt to extract spatial concepts:
@@ Appendix B User Study Details
Appendix C Extra Ablations on Two-stage Inverted Training Strategy
@@
We provide additional quantitative ablation results under the same settings in §4.3 to facilitate the understanding of the two-stage inverted training strategy. Results are shown in Tab. 3, where Two-stage means that training the global binder before training the whole hierarchical binder structure, and Inverted stands for focusing more on high noise levels in the first stage. We can observe that both techniques are crucial for achieving satisfactory optimization of the binders.
Appendix D Additional Qualitative Comparisons
@@
Figure 9: Additional Qualitative ComparisonsD). The input visual concepts and composed prompts are on the left.
@@
We provide more composed videos in Fig. 9 for additional qualitative comparisons with other methods. Fig. 9(a) demonstrates a motion transfer task, where Textual Inversion [textualinversion] and DreamVideo [dreamvideo] fails to combine the visual concepts. DualReal [wang2025dualrealadaptivejointtraining] suffers from inadequate visual concept preservation and unintended concept leakage (e.g., the green leaves). Although DB-LoRA [dreambooth] mostly follows the designated prompt to integrate visual concepts, there are significant drifts of visual concepts from the original inputs (e.g., the direction of the squirrel). BiCo achieves the best result in composing the visual concepts according to the given prompt while maintaining the consistency of visual concepts with the input image and video.
@@
Fig. 9(b) illustrates a creative style transfer task to integrate the line art sketch style with the subject in a video. All previous methods [textualinversion, dreambooth, dreamvideo, wang2025dualrealadaptivejointtraining] fail in this task to learn and compose the style concept. This sample further verifies the flexible versatile controllability of BiCo.
Appendix E Extra Case Study
@@
Figure 10: Extra Case Study for ComponentsE). The input visual concepts and composed prompts are on the left.
@@
We further illustrate the functions of BiCo’s components with another concrete visual concept composition sample in Fig. 10. Comparing #2 to #1, we can observe that the hierarchical binder structure enables our method to encode more visual information into binders, resulting in better concept preservation results. The prompt diversification operation (#3) and the absorbent token (#4) in DAM enhance the accuracy of concept-prompt binding, better preserving background details in the composed videos. The effectiveness of the absorbent token can also be verified by the enhanced background preservation in #7 compared to #5. TDS further improves the composition quality by enhancing the compatibility between image and video concepts, as illustrated by comparing #7 to #4 and #5 to #3. The two-stage inverted training strategy significantly stabilizes the optimization process, bringing considerably better results in the same optimization steps (#7 to #6). The video results can be found in the supplemental page.
Appendix F More Discussions
F.1 Limitations
@@
Figure 11: Failure CasesF.1). In each case, the upper row shows the visual inputs, and the lower row presents the composed video.

### v2

@@ 4.2.1 Quantitative Comparisons
We construct 40 test cases with images and videos from the DAVIS [43] dataset and the Internet for evaluation. Both automatic metrics and human evaluations are adopted for assessing the concept composition performance. For automatic metrics, we use CLIP-T to measure the alignment between the generated video and the textual prompt with CLIP [45] feature similarities, and choose DINO-I to quantify the preservation of visual concepts with the harmonic mean of DINO [10] feature similarities between the composed video and all visual inputs. For human evaluations, we asked 28 volunteers to rate the composed video in the following aspects with a 5-point Likert scale: 1) Concept Preservation: how well the composed video preserves the concepts from corresponding visual sources. 2) Prompt Fidelity: how well the composed video follows the input prompt. 3) Motion Quality: the motion quality of the composed video considering motion smoothness, consistency, naturalness, etc. We compute the average score of the 3 aspects as the Overall Quality. Please refer to the supplementary for more details on user study settings.
Table 1: Quantitative Comparisons with Prior Arts4.2.1). Results in bold are the best. † Implemented on Wan2.1 [54].
@@
As displayed in Tab. 1, BiCo consistently outperforms all other methods in both automatic metrics and human evaluations. Compared to the prior art DualReal [55], our method achieves a +54.67% improvement on the subjective Overall Quality. In addition, BiCo also supports the extraction of non-object concepts, learning multiple concepts from a single input, arbitrary image/video input types, and flexible concept composition via prompt manipulation, where previous methods fall short.
@@ 4.2.2 Qualitative Comparisons
We visualize the composed videos in Fig. 5 to provide an intuitive comparison with other methods. It shows a creative motion transfer task, where Textual Inversion [17] and DreamVideo [56] fails to combine the visual concepts. DualReal [55] does not accurately follow the prompt to compose the concepts, and the generated video is almost static. Although DB-LoRA [48] mostly follows the designated prompt to integrate visual concepts, there are significant drifts of visual concepts from the original inputs. BiCo best composes the visual concepts according to the given prompt while maintaining the visual concept consistency with the input image and video.
@@ 5 Conclusion and Discussion
Acknowledgment
This work was partially supported by a grant from the NSFC/RGC Collaborative Research Scheme Project No. CRS_HKUST605/25.
@@ Acknowledgment
  • §C provides more details on the two-stage inverted training strategy and conducts further ablations.
  • §D explains the justifications for the absorbent token and provides empirical evidence.
  • §E illustrates more qualitative comparisons.
  • §F performs another case study to facilitate the understanding of different components of BiCo.
  • §G further discusses the limitations with failure cases and the societal impacts of BiCo.
Please refer to the webpage for video results.
@@ Appendix A Detailed Prompts for DAM
In the prompt diversification process, we utilize a powerful VLM Qwen2.5-VL [7] to generate diversified concept prompts while retaining the key conceptual words unchanged. During the key concept extraction stage, the VLM is asked to extract essential spatial and temporal concepts from the visual inputs. For image inputs, we use the following textual prompt to extract spatial concepts:
@@ Appendix B User Study Details
Appendix C Extra Details and Ablations on Two-stage Inverted Training Strategy
The probability distribution for the discretized timestep $t_{i}$ in inverted training is:
$p(t_{i})=\begin{cases}(1-\beta)\cdot\frac{1}{N_{<\alpha}}&,d(t_{i})<\alpha\\ \beta\cdot\frac{1}{N_{\geq\alpha}}&,d(t_{i})\geq\alpha\end{cases},$ (4)
where $d(t_{i})\in[0,1]$ indicates the position of $t_{i}$ in the scheduler, and $N_{*}$ is the total number of discretized timesteps in the interval *. $\alpha=0.875$ is selected according to the training recipe of Wan2.2 to distinguish higher and lower noise levels. While $\beta$ can be selected in a reasonable range to emphasize the higher noise levels, we empirically found that setting $\beta=\alpha$ exchanges the total probability mass between the higher and lower noise levels and yields satisfactory performance given that higher noise levels originally account for a smaller probability than lower noise levels.
@@ Appendix C Extra Details and Ablations on Two-stage Inverted Training Strategy
We provide additional quantitative ablation results under the same settings in §4.3 to facilitate understanding of the two-stage inverted training strategy. Results are shown in Tab. 3, where Two-stage means that training the global binder before training the whole hierarchical binder structure, and Inverted stands for focusing more on high noise levels in the first stage. We can observe that both techniques are crucial for achieving satisfactory optimization of the binders.
Appendix D Analysis on Absorbent Token
In T2V models, text tokens are already associated with corresponding visual concepts as a good initial value for further personalization. This association is the foundation for our binders to learn sample-specific features. With a new absorbent token, it is expected that the model encodes irrelevant information into this token instead of other tokens with good initialization for corresponding visual concepts. The absorbent token is expected not to capture specific concepts, but to prevent other conceptual tokens from being distracted from established initial associations.
We demonstrate the effectiveness of the absorbent token by reconstructing a target image with the trained binder and visualizing the cross-attention maps of the target subject tokens (Akita dog) and the absorbent token in Fig. 9. As observed, the absorbent token does capture irrelevant details like plants. Removing the trained absorbent token during inference also enhances attention on the target.
Figure 9: Visualizations of cross-attention maps of target subject tokens (Akita dog) (§D).
Appendix E Additional Qualitative Comparisons
@@
Figure 10: Additional Qualitative ComparisonsE). The input visual concepts and composed prompts are on the left.
@@
We provide more composed videos in Fig. 10 for additional qualitative comparisons with other methods. Fig. 10(a) demonstrates a motion transfer task, where Textual Inversion [17] and DreamVideo [56] fails to combine the visual concepts. DualReal [55] suffers from inadequate visual concept preservation and unintended concept leakage (e.g., the green leaves). Although DB-LoRA [48] mostly follows the designated prompt to integrate visual concepts, there are significant drifts of visual concepts from the original inputs (e.g., the direction of the squirrel). BiCo achieves the best result in composing the visual concepts according to the given prompt while maintaining the consistency of visual concepts with the input image and video.
@@
Fig. 10(b) illustrates a creative style transfer task to integrate the line art sketch style with the subject in a video. All previous methods [17, 48, 56, 55] fail in this task to learn and compose the style concept. This sample further verifies the flexible versatile controllability of BiCo.
Appendix F Extra Case Study
@@
Figure 11: Extra Case Study for ComponentsF). The input visual concepts and composed prompts are on the left.
@@
We further illustrate the functions of BiCo’s components with another concrete visual concept composition sample in Fig. 11. Comparing #2 to #1, we can observe that the hierarchical binder structure enables our method to encode more visual information into binders, resulting in better concept preservation results. The prompt diversification operation (#3) and the absorbent token (#4) in DAM enhance the accuracy of concept-prompt binding, better preserving background details in the composed videos. The effectiveness of the absorbent token can also be verified by the enhanced background preservation in #7 compared to #5. TDS further improves the composition quality by enhancing the compatibility between image and video concepts, as illustrated by comparing #7 to #4 and #5 to #3. The two-stage inverted training strategy significantly stabilizes the optimization process, bringing considerably better results in the same optimization steps (#7 to #6). The video results can be found in the webpage.
Appendix G More Discussions
G.1 Limitations
@@
Figure 12: Failure CasesG.1). In each case, the upper row shows the visual inputs, and the lower row presents the composed video.